HRMS, 28th Colloquium, Dijon, France, Aug. 28 - Sept. 1, 2023

MODIFIED COMPLEX ROBERT-BONAMY LINE SHAPE CALCULATIONS FOR THE CO_{2}-H_{2}O COLLISION SYSTEM

\underline{R. GAMACHE} , University of Massachusetts Lowell, 265 Riverside Street, 01854, Lowell, MA, USA; \underline{B. VISPOEL} , Research unit Lasers and Spectroscopies (LLS), Institute of Life, Earth and Environment (ILEE), University of Namur, 61 Rue de

Bruxelles, Namur, Belgium

Modified Complex Robert-Bonamy (MCRB) calculations were made for the CO_{2}- H_{2}O collision system using an intermolecular potential comprised of quadrupoledipole, quadrupole-quadrupole, atom-atom component expanded to 20^{th} order and rank 4, induction and London dispersion terms. The intermolecular potential was refined using the measurements of Sung et al.^{1} for 46 ν_{3} transitions up to J”=60. The final potential gives results that compare very well with the measurements of Sung et al.; average percent difference of 0.07 and a standard deviation of 1.10 percent. See the figure below.

This intermolecular potential was used to make calculations for all transitions in the (00001)-(00011), (00001)-(00001), and (01101)-(00001) bands for J up to 120 and a temperature range of 200-3000 K. The temperature dependence of the halfwidths and line shifts were determined via the Gamache-Vispoel model^{2}. Calculations are being made for other bands with large numbers of ν_{1}, ν_{2}, and ν_{3} quanta exchanged in order to study the vibrational dependence of the half-widths and line shifts and to develop a prediction routine.^{3}.

1doi:10.1139/P08-130, K. Sung, L.R. Brown, R.A. Toth, T.J. Crawford, Can. J. Phys. 87, 469-484, (2009).

2doi:10.1016/j.jqsrt.2018.05.019, R.R. Gamache and B. Vispoel, J. Quant. Spectrosc. Radiat.

Transfer 217, 440-452, (2018).

3doi:10.1016/j.jqsrt.2013.05.021, R.R. Gamache and J. Lamouroux, J. Quant. Spectrosc. Radiat.

Transfer 117, 93-103, (2013).

p-number: p034  Submitted on Mon, 05 Jun 2023 14:47:24 +0200